# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v1

@@ 5 Discussion and Conclusion
In this paper, we introduce InvestESG, a MARL environment simulating long-term interactions between companies and investors under varying ESG disclosure policies. For policymakers, InvestESG shows that mandatory ESG disclosure, paired with informed, ESG-conscious investors, can drive corporate mitigation efforts. Additionally, providing high-quality system-wide information effectively motivates action from both corporations and investors. For economics and policy researchers, InvestESG introduces MARL as a powerful tool to complement traditional empirical and theoretical methods, allowing scalable policy testing in a simulated environment. Our model predicts agent behaviors consistent with empirical evidence and uncovers novel insights. For example, greenwashing poses a lesser challenge than anticipated, and providing system-level climate risk information helps resolve the dilemma even without investor involvement. These findings position InvestESG as a promising, ecologically valid platform for testing more complex and nuanced policies. For the machine learning community, InvestESG presents a novel multi-agent benchmark, fostering the development of RL algorithms that tackle complex social dilemmas, competition, and long-term strategy—pushing forward AI applications in real-world, high-impact domains. We encourage the machine learning community to develop algorithmic innovations for InvestESG that can inspire real-world actions to address climate change.
Limitations. We acknowledge that InvestESG simplifies various aspects of real-world climate evolution, financial markets, and regulatory frameworks; for example, the cost of mitigation efforts varies by market sector and geographic location, and investor ESG preferences can fluctuate with market conditions and social trends. These simplifications were intentional to enable flexible experimentation on the core dynamics between companies and investors at this stage. However, we anticipate ongoing improvements, as outlined below.
@@
Future Work. Our goal is to spark discussions and evolve InvestESG into a robust tool for policy design, with input from experts in machine learning, climate change, and economics. We envision three key areas for future work: (1) Richer environment. For instance, we plan to incorporate employees and customers, who may also be attracted to more climate-friendly companies. We also envision capturing complex market and environmental dynamics between industries, and enhancing investor decision-making to move beyond binary choices to better reflect real-world complexities. (2) Enhanced policy design. We propose testing policies that strengthen ESG disclosure mandates, such as guidelines on Scope 1 and 2 emissions, which, while costly, may ease the evaluation of companies’ efforts. We also suggest exploring ways to redesign the simple ESG score to better inform investors and motivate corporate actions. (3) Private-sector initiatives. Given the challenges of passing new regulations, we aim to explore if companies and investors can self-regulate through coordination. For instance, we suggest testing if collective funds or peer monitoring can enhance mitigation or lead to collusion. We invite researchers interested in multi-agent learning to collaborate with us on advancing these ideas and turning InvestESG into a comprehensive policy tool.
@@ 7 Reproducibility Statement
To promote reproducibility, we provide our source code and corresponding configuration files used for running the experiments as supplementary material. The details about installation and sample code for running the experiment is also included in the supplementary material. To encourage engagement and exploration from the machine learning community, we provide open-source code for the InvestESG environment and baselines via Github.

### v2

@@ 5 Main Results
In addition to the experiments presented above, we also explore settings where (1) company agents are allowed to spend on building their own resilience, (2) company agents are initialized with real-world company actions, (3) agents’ decisions are locked in for five years to simulate capital flexibility challenges, (4) companies’ climate resilience parameter $L_{t}^{\mathcal{C}_{i}}$ are random and vary across events and companies, and (5) a more strict bankruptcy mechanism. These experiments yield directional conclusions consistent with those presented in the main text. Detailed results can be found in Appendix 11.
@@ 6 Discussion and Conclusion
In this paper, we introduce InvestESG, a MARL environment simulating long-term interactions between companies and investors under varying ESG disclosure policies. For policymakers, InvestESG shows that mandatory ESG disclosure, paired with highly ESG-conscious investors, can drive corporate mitigation efforts. Additionally, greenwashing poses a lesser challenge than anticipated, and providing high-quality system-wide information effectively motivates action from both corporations and investors. For economics and policy researchers, InvestESG introduces MARL as a promising tool to complement traditional empirical and theoretical methods, allowing scalable policy testing in a simulated environment. Our model predicts agent behaviors consistent with empirical evidence and uncovers novel insights. For the machine learning community, InvestESG presents a novel multi-agent benchmark, fostering the development of RL algorithms that tackle complex social dilemmas, competition, and long-term strategy—pushing forward AI applications in real-world, high-impact domains. We encourage the machine learning community to develop algorithmic innovations for InvestESG that can inspire practical actions to address climate change.
Limitations. We acknowledge that InvestESG simplifies various aspects of real-world climate evolution, financial markets, and regulatory frameworks. These simplifications were intentional for two reasons: (1) rather than fully simulating the global market and climate complexities, which is arguably infeasible for a single model, we aim to develop a "first-principles" model that offers directional insights for policy design; (2) we designed the benchmark to be simple enough for the ML community to iterate on with reasonable computational resources (e.g., a single GPU). By lowering computational barriers, we hope to encourage broader participation from the ML community.
At this stage, InvestESG effectively captures the core incentive structures necessary to evaluate the policy in question. We anticipate future improvements to include more complexity, allowing researchers and policymakers to select the appropriate level of granularity for their needs.
@@
Future Work. Our goal is to spark discussions and evolve InvestESG into a robust tool for policy design, with input from experts in machine learning, climate change, and economics. We envision three key areas for future work: (1) Richer environment. For instance, we plan to incorporate employees and customers, who may also be attracted to more climate-friendly companies. We also envision capturing complex market and environmental dynamics between industries (e.g., include insurance companies (Brogi et al., 2022)), and enhancing investor decision-making to move beyond binary choices to better reflect real-world complexities. (2) Enhanced policy design. We propose testing policies that strengthen ESG disclosure mandates, such as guidelines on Scope 1 and 2 emissions, which, while costly, may ease the evaluation of companies’ efforts. We also suggest exploring ways to redesign the simple ESG score to better inform investors and motivate corporate actions. (3) Richer agent dynamics. For instance, we plan to allow investors to learn ESG-consciousness level through interaction with the environment, which can change overtime. Given the challenges of passing new regulations, we aim to explore if companies and investors can self-regulate through coordination. For instance, we suggest testing if collective funds or peer monitoring can enhance mitigation or lead to collusion.
@@ 8 Reproducibility Statement
To promote reproducibility, we provide our source code and corresponding configuration files used for running the experiments as supplementary material. The details about installation and sample code for running the experiment is also included in the supplementary material. To encourage engagement and exploration from the machine learning community, we will open source the code for the InvestESG environment and baselines via Github.